Merone Company allocates materials handling cost to the company's two products using the below data:
Modular Homes   Prefab Barns
Total expected units produced   6,100   9,100
Total expected material moves   610   210
Expected direct labor-hours per unit   810   310
The total materials handling cost for the year is expected to be $388,100.
If the materials handling cost is allocated on the basis of material moves, the total materials handling cost allocated to the modular homes is closest to: (Round your intermediate calculations to 2 decimal places.)
Multiple Choice
$288,706.90
$280,679.46
$196,123.00
$262,906.00
Answer #1

Answer

· Total Moves = 610 + 210 = 820

· Activity rate = $ 388100 / 820 moves = $ 473.29

· Material handling cost allocated to Modular homes
= 610 moves x $ 473.29
= $ 288,706.90

· Correct Answer = Option #1: $ 288,706.90